Consider the system of equations

11 2x1 + x2 + 4x3 =

3x1 + x2 + 5x3 =14

A feasible solution is x1 = ,2 x2 = ,3 x3 = .1 Reduce this feasible solution to a basic

feasible solution.



1
Expert's answer
2021-04-28T04:09:08-0400

"\begin{cases}
 2x_1+x_2=11-4x_3, \\
 3x_1+x_2=14-5x_3;
\end{cases}"


"\Delta=\begin{vmatrix}
 2 & 1 \\
 3 & 1
\end{vmatrix}=2-3=-1,"


"\Delta_1=\begin{vmatrix}
 11-4x_3& 1 \\
 14-5x_3& 1
\end{vmatrix}=11-4x_3-14+5x_3=x_3-3,"


"\Delta_2=\begin{vmatrix}
 2 & 11-4x_3 \\
 3 & 14-5x_3
\end{vmatrix}=28-10x_3-33+12x_3=2x_3-5,"


"x_1=\frac{\Delta_1}{\Delta}=3-x_3,"

"x_2=\frac{\Delta_2}{\Delta}=5-2x_3."